Sally travels 100m west 240m south find the displacement and distance?
1
Expert's answer
2020-03-10T11:17:04-0400

Distance:


"d=100+240=340\ m"


Displacement:


"s=\sqrt{100^2+240^2}=260\ m"

Direction:


"\theta=\arctan{\frac{100}{240}}=22.6\degree\text{W of S}"